Convert 1228 kg/ML into g/L

Significant figures:

2.54 cm = 1 in

12 = 1 dozen

89.3×77.0×0.08

(5.01×105)÷(3.78×102)

4.005×74×0.0087

453÷2.031
Answer #1

convert 1228 kg/ML into g/L:

1228 kg/ML = 1228 * 1000 g / 0.001 L = 1.228 * 10^9 g / L

1.228 * 10^9 g / L have four significant figures.

89.3×77.0×0.08 = 550. (answer).....> three significant figures.

(5.01×105)÷(3.78×102) = 526 ÷ 385 = 1.36 (answer)........> three significant figures.

4.005×74×0.0087 = 2.6 (answer)....two significant figures.

453÷2.031 = 223. (answer) .................two significant figures.